# Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning

### v2

@@ Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea.
Appendix A Instruction Templates.
We use text templates for input texts of the utilized datasets. Specifically, the text templates consist of an instruction and an input text, as shown in Table 3. For the prediction, we utilize the representation of the mask token (i.e., "<mask>") to utilize the pre-trained knowledge of encoder-only language models.
Appendix B Module-specific Task-localization Details
@@
We conduct experiments for various modules in the transformer architecture, as described in Section 5.4. We implement algorithms for pruning those networks following our Skeleton framework to satisfy the dimension of adjacent networks after pruning. For example, suppose that we prune the Key network, $W^{K}\in\mathbb{R}^{d\times d}$, and get a pruned weight, $\tilde{W}^{K}\in\mathbb{R}^{d\times d^{\prime}}$, in the Attention module. Then, we should also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since they are adjacent networks. Therefore, after the pruning process, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by adding zero values to ensure the identical dimension. We also implement other layers in the same way with it.

### v3

@@ Acknowledgements.
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea. The Institute of Engineering Research at Seoul National University provided research facilities for this work.
GenAI Usage Disclosure
Generative AI was not used at any stage of the research, except for minor corrections related to spelling or grammar in the writing.
Appendix
Appendix A Module-specific Task-localization Details
@@
We conduct experiments for various modules in the transformer architecture, as described in Section 5.4. We implement algorithms for pruning networks following our DoE framework to satisfy the dimension of adjacent networks after pruning. For example, suppose that we prune the Key network, $W^{K}\in\mathbb{R}^{d\times d}$, and get a pruned weight, $\tilde{W}^{K}\in\mathbb{R}^{d\times d^{\prime}}$, in the Attention module. Then, we also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since it is an adjacent network. Therefore, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks after pruning, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by inserting zero values to ensure the identical dimension. We also implement other layers in the same way.